# Adaptive Agents in Spatial Double-Auction Markets: Modeling the Emergence of Industrial Symbiosis

### v1

@@ Abstract.
Agent Based Simulation, Industrial Symbiosis, Circular Economy, Market Analysis
@@ 1. Introduction
Despite its potential, IS rarely emerges spontaneously. Exchanges depend on local supply-demand compatibilities, economic incentives, and logistical constraints henriques2021industrial. Poor coordination often results in wasted resources and missed opportunities, highlighting the gap between individual strategies and collective circularity goals. Identifying the conditions under which decentralized firms can coordinate effectively is therefore a key research challenge.
@@
Analytical solutions are often intractable due to the combinatorial complexity of interactions and the emergent nature of prices and traded quantities. Simulation-based approaches, by contrast, are particularly suited to study such nonlinear, decentralized, and heterogeneous systems he2019heterogeneous. By systematically varying system configurations, simulations can indeed reveal market behaviors and highlight factors that support efficient and resilient local circularity. Agent-based models, in particular, are able to capture heterogeneity among actors and test how different market designs, decision rules, and learning mechanisms affect the emergence of exchanges under realistic economic constraints epstein1996growing. Their ability to explicitly represent geosituated agents makes them especially suitable for studying spatially dependent interactions, such as proximity based collaborations, transportation costs, and localized resource flows. Unlike theoretical game-theoretic models, which often rely on a small number of representative players to ensure tractability, agent-based models can accommodate many heterogeneous actors and interactions.
However, their effectiveness is often constrained by the simplifications made in representing agent behaviors. In many existing symbiosis models Yazan2020; Ghali2017; Raimbault2020; Gan2015; FernandezMena2020; Chahla2019, firms follow fixed behavioral rules or static pricing strategies, limiting their ability to capture the adaptive and uncertain nature of real-world exchanges. In practice, firms continuously adjust decisions in response to shifting demand, fluctuating costs, and evolving environmental regulations. This highlights the need to enhance decision-making mechanisms within agent-based models, enabling agents to learn, adapt, and explore strategies in a more dynamic way.
@@
In this paper, we present a spatially explicit, decentralized multi-agent model in which buyers and sellers of byproducts interact through multilateral spatial double auctions using co-evolving strategies. Sellers adaptively update their pricing strategies via reinforcement learning in a partially observable environment, while buyers select offers that maximize utility subject to spatial and economic constraints. A key feature of the model is that both transaction prices and traded quantities emerge endogenously from these interactions, allowing us to investigate how local decisions give rise to systemic patterns. Our contributions are threefold. First, we develop a reinforcement learning-based market model for decentralized byproduct exchanges, based on a novel auction mechanism adapted to spatial and economic constraints of territorial circularity. Second, we validate the learned strategies via counterfactual regret analysis, showing convergence toward near-Nash equilibria. Finally, a simulation-based analysis of conditions under which adaptive pricing and agent interactions lead to stable, efficient, and circular outcomes, providing insights for policymakers.
@@ 2.1. Industrial Symbiosis and Circular Economy Modeling
Industrial symbiosis in Eco-Industrial Parks (EIPs) has traditionally been studied using system dynamics to capture network level behavior and feedback loops zhao2008simulation; qu2010system; sterman2000business; lorenz2006towards; cao2009applying. Complementary approaches have treated EIPs as complex adaptive systems, highlighting dynamic business interactions and emergent symbiotic relationships chertow2012organizing; yu2013process; felicio2016; lambert2002eco; romero2013framework. Other studies focused on material and energy flows, byproduct exchanges, and social network analysis Romero2014; felicio2016; yazan2016design; chertow2000; domenech2009social. Recent reviews Demartini2022; rizzati2024systematic emphasize the potential of agent-based modeling to capture local, adaptive, and self-organizing behaviors. Emerging agent-based studies have integrated input-output models Yazan2020, explored spatially networked exchanges Ghali2017; Raimbault2020; mastioCIEC, and examined sector specific synergies Gan2015; FernandezMena2020; Chahla2019. Yet, despite these advances, the explicit representation of economic incentives and market coordination mechanisms remains underexplored. Also, these works does not explicitly model markets where prices, trade volumes, and resource allocations emerge endogenously from decentralized interactions.
@@ 2.2. Auctions and Market Mechanisms in Spatial Settings
Auction and negotiation mechanisms in Multi-Agent Systems (MAS) have become canonical tools for decentralized coordination. From early negotiation platforms Chavez1997 to double auctions—market mechanisms in which multiple buyers and sellers simultaneously submit bids and offers—used to study bidding strategies and price discovery Tesauro2001; Satterthwaite2022, research has demonstrated how decentralized negotiations can generate robust outcomes. Extensions have examined repeated competition Ledvina2010; Bichler2024, peer-to-peer trading Qiu2021, social influence wainwright2024social, and multilateral bargaining across multiple resources Chen2015; An2011. Spatial aspects are frequently acknowledged, for instance through costly trade links babaioff2005, heterogeneous bilateral transport costs chu2008, or land markets where localized interactions induce differentiated prices filatova2007. Yet these contributions remain largely theoretical, typically involving few participants and limited attention to adaptive mechanisms that capture the evolving dynamics of real markets. In particular, the integration of spatial frictions with learning-based bidding strategies has received limited attention, despite its relevance for understanding decentralized coordination in resource-constrained environments.
@@ 2.3. Learning Agents for Adaptive Market Participation
Traditional learning agents often rely on fixed rules Cliff2024, or supervised training lauretto2013evaluation, limiting their ability to adapt to dynamic market conditions and strategic interactions. Reinforcement Learning (RL) overcomes these limitations by allowing agents to learn optimal policies through trial-and-error, balancing exploration and exploitation in complex, evolving environments. In the context of double auctions, RL agents are particularly well suited due to the complex interactions between buyers and sellers and the continuous feedback provided by order book dynamics. Agents can be designed to adopt specific roles, such as deciding when to buy, sell, or set prices, enabling cooperative and strategic trading behaviors bjerkoey2019; maeda2020. Multi-Agent Reinforcement Learning (MARL) extends this approach to settings with multiple interacting agents, allowing the simulation of bottom-up market dynamics shavandi2022; karpe2020; lussange2020; ganesh2019. MARL models have been shown to reproduce emergent behaviors in order books and centralized market microstructures by capturing interactions among heterogeneous agents karpe2020. This enables detailed analysis of price formation, liquidity, and strategic adaptation. However, existing approaches largely ignore spatial constraints, as these are typically absent in the stock market. Yet, such constraints can be critical in markets where transportation or delivery costs influence trading decisions rodrigue2020geography. RL with spatial considerations have been primarily explored in logistics, where carriers and shippers act as bidders in dynamic auctions mediated by centralized brokers Wang2018; Qiao2019; zha2017surge; atasoy2020. Even when shippers themselves are modeled as learning agents VanHeeswijk2019, exchanges remain intermediary mediated. In contrast, many real-world IS transactions are inherently local and broker-free. To the best of our knowledge, RL has not yet been applied to decentralized, spatially constrained markets where transport costs and policy incentives influence agents’ strategies.
@@ 2.4. Positioning of This Work
    Industrial symbiosis models rarely provide in depth mechanism for endogenous market-based price formation.
@@
    Reinforcement learning in markets has not been applied to decentralized settings where spatial topology influence agents behavior.
Our contribution is to bridge these strands by introducing a decentralized auction model for industrial symbiosis where (i) sellers adapt prices through reinforcement learning, (ii) spatial constraints are explicitly internalized, and (iii) both prices and traded quantities emerge from multilateral interactions. This framework connects industrial symbiosis modeling with mechanism design and multi-agent RL, allowing modelers to explore how local adaptive dynamics shape circularity outcomes. Our objective is to reproduce the systemic complexity of inter-firm interactions within industrial symbiosis as a subset of the broader circular economy, providing a foundational basis for modeling circular economic dynamics.
@@ 3.1. Environment and Agents
To concentrate on the essential dynamics of the model, we restrict our analysis to the exchange of a single byproduct. This simplification enables us to isolate and assess the effects of multilateral interactions and sellers strategies, without additional complexity introduced by multiple interdependent byproducts. We model a decentralized market for one byproduct, composed of buyers $\mathcal{B}=\{B_{1},\dots,B_{N_{B}}\}$ and sellers $\mathcal{S}=\{S_{1},\dots,S_{N_{S}}\}$. Each agent has a spatial location, and the distance between a buyer $B_{i}$ and a seller $S_{j}$ is denoted $d_{ij}$. Buyers have heterogeneous initial demands $q_{i}$ drawn from a scaled uniform distribution and evaluate incoming offers based on their price sensitivity $\beta_{i}$. Specifically, buyer $B_{i}$ accepts an offer $p_{ij}$ from seller $S_{j}$ only if $p_{ij}\leq\beta_{i}\,p_{m}$, where $p_{m}$ is the reference market price. This mechanism ensures that buyers only engage in transactions that meet their cost expectations relative to the external market.
@@
The term $c_{ij}$ is originally defined as $c_{ij}=d_{ij}\ c_{t}$, where $c_{t}$ represents the transportation cost per kilometer (expressed in percentage of the market price). This term can be further extended to capture the total logistics cost required for the byproduct provided by the seller to become usable by the buyer, including handling, treatment, or adaptation costs. Finally, $\phi_{j}$ is an adaptive pricing parameter, that evolves during the learning phase.
In addition to market and transportation costs, sellers face a penalty $c_{d}$ for unsold quantities (also expressed in percentage of the market price), representing the cost of sending residual flows to landfill and reflecting both economic and environmental impacts. If a seller $j$ ends a timestep with $q_{j}^{\text{unsold}}$ units remaining, its reward is reduced by $q_{j}^{\text{unsold}}\ c_{d}$, imposing a direct financial penalty that encourages sellers to adjust pricing strategies to minimize waste while still aiming to maximize profit.
@@
This measure allows us to systematically investigate the effects of supply-demand imbalances on agents’ behavior in the decentralized market.
Each seller $S_{j}$ aims to maximize its cumulative expected profit at each simulation step. Let $\mathbf{q}_{j}^{t}=(q_{ij}^{t})_{i\in\mathcal{B}}$ denote the vector of quantities bought by $B_{i}$ from $S_{j}$ at step $t$, and $\mathbf{p}_{j}^{t}=(p_{ij}^{t})_{i\in\mathcal{B}}$ the corresponding price vector. Defining $\mathbf{c}_{j}=(c_{ij})_{i\in\mathcal{B}}$ as the transport cost vector, the instantaneous profit can be written compactly as
@@
where $\langle\cdot,\cdot\rangle$ denotes the dot product. The first term corresponds to the net revenue from executed contracts, while the second term penalizes unsold inventory through the landfill cost $c_{d}$. This spatially defined market closely resembles a spatial Bertrand oligopoly hotelling_stability_1929, where multiple sellers compete by setting prices and buyers prefer lower-cost options. Unlike the classical Bertrand model, where identical products drive prices toward marginal cost, the inclusion of transport costs differentiates sellers spatially, allowing local price variation and creating room for adaptive strategies to emerge.
@@ 3.2. Decentralized Multilateral Auction
We model the interactions between buyers and sellers using a decentralized multilateral auction described in Algorithm 1. Each seller submits a set of bids to each buyers, where each bid specifies a price and quantity for a given product $p$, computed on the basis of the logistic costs required to deliver the product to the buyer’s location and the external market price. Buyers evaluate the incoming bids and select the one that maximizes their utility, subject to their price sensitivity $\beta_{i}$ and demand $q_{i}$. Selected bids are then sent back as proposals to sellers. Each seller chooses a subset of the received bids to accept in order to maximize expected profit, taking into account transport costs $c_{ij}$ and remaining inventory $q_{j}$. Contracts are executed, updating both buyers’ needs and sellers’ available quantities. This process iterates until no further mutually acceptable bids can be matched. Sellers adapt their pricing strategy $\phi_{s}$ over time using a strategy based on historical rewards that will be described in 3.3.
@@
2:while contracts has been made at the previous timestep do
@@
4:  Buyer choice: each $B_{i}$ evaluates the received bids, selects the most attractive one within price tolerance $\beta_{i}p_{m}$, and sends acceptance proposal to the corresponding seller
@@ 3.3. Seller Reinforcement Learning and Action Selection
Here, $\alpha\in[0,1]$ is the smoothing factor, which balances the influence of past and recent observations. By prioritizing recent data, EMA allows rapid adaptation to changing conditions while remaining computationally efficient, requiring only constant-time updates per action. Although it does not offer formal convergence guarantees, EMA provides a practical compromise between adaptivity and scalability, and allows to maintain high learning efficiency morales2024exponential.
Sellers select their next action using Boltzmann (softmax) sampling sutton1998reinforcement with a declining temperature $\tau_{t}$:
@@
where $\tau_{0}$ sets the initial exploration level, decay controls the annealing rate, and $\tau_{\min}$ prevents premature convergence. This ensures exploration dominates early in learning, while the policy gradually becomes greedier over time. The subsequent action $\phi_{j}^{(t+1)}$ is drawn at random from the set of actions according to probabilities $P_{k}$. This sampling strategy balances exploration and exploitation in a multi-agent, stochastic environment. In our simulations, Boltzmann exploration consistently converged faster and with less oscillation than $\epsilon$-greedy or UCB, which often struggled with noise from multi-agent interactions o2021variational. By combining a discretized action space, adaptive reward estimates, and Boltzmann sampling, the model captures the decentralized emergence of price formation and allocation dynamics in a spatially distributed market for byproducts. While Boltzmann does not always guarantee optimal regret bounds in multi-agent stochastic settings, its smooth adaptation makes it practical and effective for this scenario.

### v2

@@ Abstract.
Spatial Agent-Based Simulation, Industrial Symbiosis, Circular Economy, Distributed Market Analysis, Reinforcement Learning
doi:
EXII2056
@@ 1. Introduction
Despite its potential, IS rarely emerges spontaneously. Exchanges depend on local supply-demand compatibilities, economic incentives, and logistical constraints henriques2021industrial. Poor coordination often results in wasted resources and missed opportunities, highlighting the gap between individual strategies and collective circularity goals. Identifying the conditions under which decentralized firms can coordinate effectively is, therefore, a key research challenge.
@@
Analytical solutions are often intractable due to the combinatorial complexity of interactions and the emergent nature of prices and traded quantities. Simulation-based approaches, by contrast, are particularly suited to study such nonlinear, decentralized, and heterogeneous systems he2019heterogeneous. By systematically varying system configurations, simulations can indeed reveal market behaviors and highlight factors that support efficient and resilient local circularity. Agent-based models, in particular, are able to capture heterogeneity among actors and test how different market designs, decision rules, and learning mechanisms affect the emergence of exchanges under realistic economic constraints epstein1996growing. Their ability to explicitly represent geosituated agents makes them especially suitable for studying spatially dependent interactions, such as proximity-based collaborations, transportation costs, and localized resource flows. Unlike theoretical game-theoretic models, which often rely on a small number of representative players to ensure tractability, agent-based models can accommodate many heterogeneous actors and interactions. i However, their effectiveness is often constrained by the simplifications made in representing agent behaviors. In many existing symbiosis models Yazan2020; Ghali2017; Raimbault2020; Gan2015; FernandezMena2020; Chahla2019, firms follow fixed behavioral rules or static pricing strategies, limiting their ability to capture the adaptive and uncertain nature of real-world exchanges. In practice, firms continuously adjust decisions in response to shifting demand, fluctuating costs, and evolving environmental regulations. This highlights the need to enhance decision-making mechanisms within agent-based models, enabling agents to learn, adapt, and explore strategies in a more dynamic way.
@@
In this paper, we present a spatially explicit, decentralized multi-agent model in which buyers and sellers of byproducts interact through multilateral spatial double auctions using co-evolving strategies. Sellers adaptively update their pricing strategies via reinforcement learning in a partially observable environment, while buyers select offers that maximize utility subject to spatial and economic constraints. A key feature of the model is that both transaction prices and traded quantities emerge endogenously from these interactions, allowing us to investigate how local decisions give rise to systemic patterns. While illustrated here through industrial symbiosis, this framework applies broadly to circular economy settings, such as heat recovery, wastewater reuse, or biomass valorization, where decentralized agents, ranging from firms to aggregated entities like neighborhoods, must coordinate to close resource loops under transport frictions. Our contributions are threefold. First, we develop a reinforcement learning-based market model for decentralized byproduct exchanges, based on a novel auction mechanism adapted to spatial and economic constraints of territorial circularity. Second, we validate the learned strategies via counterfactual regret analysis, showing convergence toward near-Nash equilibrium. Finally, we propose a simulation-based analysis of the conditions under which adaptive pricing and agent interactions lead to stable circular outcomes, providing insights for policymakers.
@@ 2.1. Industrial Symbiosis and Circular Economy Modeling
Industrial symbiosis in Eco-Industrial Parks (EIPs) has traditionally been studied using system dynamics to capture network level behavior and feedback loops zhao2008simulation; qu2010system; sterman2000business; lorenz2006towards; cao2009applying. Complementary approaches have treated EIPs as complex adaptive systems, highlighting dynamic business interactions and emergent symbiotic relationships chertow2012organizing; yu2013process; felicio2016; lambert2002eco; romero2013framework. Other studies focused on material and energy flows, byproduct exchanges, and social network analysis Romero2014; felicio2016; yazan2016design; chertow2000; domenech2009social. Recent reviews Demartini2022; rizzati2024systematic emphasize the potential of agent-based modeling to capture local, adaptive, and self-organizing behaviors. Emerging agent-based studies have integrated input-output models Yazan2020, explored spatially networked exchanges Ghali2017; Raimbault2020; mastioCIEC, and examined sector specific synergies Gan2015; FernandezMena2020; Chahla2019. Yet, despite these advances, the explicit representation of economic incentives and market coordination mechanisms remains underexplored. Additionally, these works do not explicitly model markets where prices, trade volumes, and resource allocations emerge endogenously from decentralized interactions.
@@ 2.2. Auctions and Market Mechanisms in Spatial Settings
Auction and negotiation mechanisms in Multi-Agent Systems (MAS) have become canonical tools for decentralized coordination. From early negotiation platforms Chavez1997 to double auctions—market mechanisms in which multiple buyers and sellers simultaneously submit bids and offers—used to study bidding strategies and price discovery Tesauro2001; Satterthwaite2022, research has demonstrated how decentralized negotiations can generate robust outcomes. Extensions have examined repeated competition Ledvina2010; Bichler2024, peer-to-peer trading Qiu2021, social influence wainwright2024social, and multilateral bargaining across multiple resources Chen2015; An2011. Spatial aspects are frequently acknowledged, for instance through costly trade links babaioff2005, heterogeneous bilateral transport costs chu2008, or land markets where localized interactions induce differentiated prices filatova2007. Yet these contributions remain largely theoretical, typically involving few participants and limited attention to adaptive mechanisms that capture the evolving dynamics of real markets.
@@ 2.3. Learning Agents for Adaptive Market Participation
Traditional learning agents often rely on fixed rules Cliff2024, or supervised training lauretto2013evaluation, limiting their ability to adapt to dynamic market conditions and strategic interactions. Reinforcement Learning (RL) overcomes these limitations by allowing agents to learn optimal policies through trial-and-error, balancing exploration and exploitation in complex, evolving environments. In the context of double auctions, RL agents are particularly well-suited due to the complex interactions between buyers and sellers and the continuous feedback provided by order book dynamics. Agents can be designed to adopt specific roles, such as deciding when to buy, sell, or set prices, enabling cooperative and strategic trading behaviors bjerkoey2019; maeda2020. Multi-Agent Reinforcement Learning (MARL) extends this approach to settings with multiple interacting agents, allowing the simulation of bottom-up market dynamics shavandi2022; karpe2020; lussange2020; ganesh2019. MARL models have been shown to reproduce emergent behaviors in order books and centralized market microstructures by capturing interactions among heterogeneous agents karpe2020. This enables detailed analysis of price formation, liquidity, and strategic adaptation. However, existing approaches largely ignore spatial constraints, as these are typically absent in the stock market. Yet, such constraints can be critical in markets where transportation or delivery costs influence trading decisions rodrigue2020geography. RL with spatial considerations has been primarily explored in logistics, where carriers and shippers act as bidders in dynamic auctions mediated by centralized brokers Wang2018; Qiao2019; zha2017surge; atasoy2020. Even when shippers themselves are modeled as learning agents VanHeeswijk2019, exchanges remain intermediary mediated. In contrast, many real-world IS transactions are inherently local and broker-free. To the best of our knowledge, RL has not yet been applied to decentralized, spatially constrained markets where transport costs and policy incentives influence agents’ strategies.
@@ 2.4. Positioning of This Work
    Industrial symbiosis models rarely provide an in-depth mechanism for endogenous market-based price formation.
@@
    Reinforcement learning in markets has not been applied to decentralized settings where spatial topology influences agents’ behavior.
Our contribution is to bridge these strands by introducing a decentralized auction model for industrial symbiosis. While RL and double auctions have been studied in isolation, their integration into a spatial multilateral double-auction, where transport costs and location directly drive strategic behavior, constitutes a novel contribution. Specifically, our model ensures that (i) sellers adapt prices through reinforcement learning, (ii) spatial constraints are explicitly internalized, and (iii) both prices and traded quantities emerge from multilateral interactions. Our objective is to reproduce the systemic complexity of inter-firm interactions within industrial symbiosis as a subset of the broader circular economy, providing a foundational basis for modeling circular economic dynamics. Beyond the specific application, the contribution lies in the rigorous methodology and implementation of this multi-agent RL framework, offering a reusable and generalizable approach for studying spatial market interactions.
@@ 3.1. Environment and Agents
To concentrate on the essential dynamics of the model, we restrict our analysis to the exchange of a single byproduct. This simplification enables us to isolate and assess the effects of multilateral interactions and sellers’ strategies, without additional complexity introduced by multiple interdependent byproducts. We model a decentralized market for one byproduct, composed of buyers $\mathcal{B}=\{B_{1},\dots,B_{N_{B}}\}$ and sellers $\mathcal{S}=\{S_{1},\dots,S_{N_{S}}\}$. Each agent has a spatial location, and the distance between a buyer $B_{i}$ and a seller $S_{j}$ is denoted $d_{ij}$. Buyers have heterogeneous initial demands $q_{i}$ drawn from a scaled uniform distribution and evaluate incoming offers based on their price sensitivity $\beta_{i}$. Specifically, buyer $B_{i}$ accepts an offer $p_{ij}$ from seller $S_{j}$ only if $p_{ij}\leq\beta_{i}\,p_{m}$, where $p_{m}$ is the reference market price. This mechanism ensures that buyers only engage in transactions that meet their cost expectations relative to the external market. We adopt here a ”Small Open Economy”GULATI2013288 assumption: we hypothesize that local industrial park internal exchanges are too small to impact the global commodity market price. Consequently, $p_{m}$ is treated as a fixed exogenous parameter.
@@
The term $c_{ij}$ is originally defined as $c_{ij}=d_{ij}\ c_{t}$, where $c_{t}$ represents the transportation cost per kilometer (expressed as a percentage of the market price). This term can be further extended to capture the total logistics cost required for the byproduct provided by the seller to become usable by the buyer, including handling, treatment, or adaptation costs. Finally, $\phi_{j}$ is an adaptive pricing parameter that evolves during the learning phase.
In addition to market and transportation costs, sellers face a penalty $c_{d}$ for unsold quantities (also expressed as a percentage of the market price), representing the cost of sending residual flows to landfill and reflecting both economic and environmental impacts. If a seller $j$ ends a timestep with $q_{j}^{\text{unsold}}$ units remaining, its reward is reduced by $q_{j}^{\text{unsold}}\ c_{d}$, imposing a direct financial penalty that encourages sellers to adjust pricing strategies to minimize waste while still aiming to maximize profit.
@@
This measure allows us to systematically investigate the effects of supply-demand imbalances on agents’ behavior in the decentralized market. Each seller $S_{j}$ aims to maximize its cumulative expected profit at each simulation step. Let $\mathbf{q}_{j}^{t}=(q_{ij}^{t})_{i\in\mathcal{B}}$ denote the vector of quantities bought by $B_{i}$ from $S_{j}$ at step $t$, and $\mathbf{p}_{j}^{t}=(p_{ij}^{t})_{i\in\mathcal{B}}$ the corresponding price vector. Defining $\mathbf{c}_{j}=(c_{ij})_{i\in\mathcal{B}}$ as the transport cost vector, the instantaneous profit can be written compactly as
@@
where $\langle\cdot,\cdot\rangle$ denotes the dot product. The first term corresponds to the net revenue from executed contracts, while the second term penalizes unsold inventory through the landfill cost $c_{d}$. This spatially defined market closely resembles a spatial Bertrand oligopoly hotelling_stability_1929, where multiple sellers compete by setting prices, and buyers prefer lower-cost options. Unlike the classical Bertrand model, where identical products drive prices toward marginal cost, the inclusion of transport costs differentiates sellers spatially, allowing local price variation and creating room for adaptive strategies to emerge.
@@ 3.2. Decentralized Multilateral Auction
We model the interactions between buyers and sellers using a decentralized multilateral auction described in Algorithm 1. Each seller submits a set of bids to each buyer, where each bid specifies a price and quantity for a given product $p$, computed on the basis of the logistic costs required to deliver the product to the buyer’s location and the external market price. Buyers evaluate the incoming bids and select the one that maximizes their utility, subject to their price sensitivity $\beta_{i}$ and demand $q_{i}$. Selected bids are then sent back as proposals to sellers. Each seller chooses a subset of the received bids to accept in order to maximize expected profit, taking into account transport costs $c_{ij}$ and remaining inventory $q_{j}$. Contracts are executed, updating both buyers’ needs and sellers’ available quantities. This process iterates until no further mutually acceptable bids can be matched. Sellers adapt their pricing strategy $\phi_{s}$ over time using a strategy based on historical rewards that will be described in 3.3.
@@
2:while contracts have been made at the previous timestep do
@@
4:  Buyer choice: each $B_{i}$ evaluates the received bids, selects the most attractive one within price tolerance $\beta_{i}p_{m}$, and sends an acceptance proposal to the corresponding seller
@@ 3.3. Seller Reinforcement Learning and Action Selection
Here, $\alpha\in[0,1]$ is the smoothing factor, which balances the influence of past and recent observations. By prioritizing recent data, EMA allows for rapid adaptation to changing conditions while remaining computationally efficient, requiring only constant-time updates per action. Although it does not offer formal convergence guarantees, EMA provides a practical compromise between adaptivity and scalability and allows for maintaining high learning efficiency morales2024exponential. Sellers select their next action using Boltzmann (softmax) sampling sutton1998reinforcement with a declining temperature $\tau_{t}$:
@@
where $\tau_{0}$ sets the initial exploration level, decay controls the annealing rate, and $\tau_{\min}$ prevents premature convergence. This ensures exploration dominates early in learning, while the policy gradually becomes more conservative over time. The subsequent action $\phi_{j}^{(t+1)}$ is drawn at random from the set of actions according to probabilities $P_{k}$. This sampling strategy balances exploration and exploitation in a multi-agent, stochastic environment. In our simulations, Boltzmann exploration consistently converged faster and with less oscillation than $\epsilon$-greedy or UCB, which often struggled with noise from multi-agent interactions o2021variational. By combining a discretized action space, adaptive reward estimates, and Boltzmann sampling, the model captures the decentralized emergence of price formation and allocation dynamics in a spatially distributed market for byproducts. While Boltzmann does not always guarantee optimal regret bounds in multi-agent stochastic settings, its smooth adaptation makes it practical and effective for this scenario.